Find the specified probability, from a table of Normal probabilities. Assume that the necessary conditions and assumptions are met. The annual precipitation amounts in a certain mountain range are normally distributed with a mean of 77 inches, and a standard deviation of 10 inches. What is the probability that the mean annual precipitation during 25 randomly picked years will be less than 79.8 inches?
Answer #1

Given,

\mu = 77 , \sigma = 10

Using central limit theorem,

P(\bar{x} < x) = P (Z < x - \mu / (\sigma / sqrt(n) ) )

So,

P( \bar{x} < 79.8 ) = P( Z < 79.8 - 77 / ( 10 / sqrt(25) ))

= P (Z < 1.4)

= 0.9192 (Probability calculated from Z table)